12 hours : 30 hours = 8km : 20km
 12 hours : 30 hours = 8km : 20km

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is True
Consider the given ratios, 12 hours : 30 hours = 8km : 20km

12/30 = 8/20
12/30 is further simplified by dividing both numerator and denominator by 6 we get,

= 2/5
8/20 is further simplified by dividing both numerator and denominator by 4 we get, = 2/5
Hence, 2/5 = 2/5
Therefore, 12 hours : 30 hours = 8km : 20km

Thanks